# Review on the decoding algorithms for surface codes

### v4

@@ 4.2 Twirled quantum channels
Therefore, the GAPD channel is a complete mathematical description of the evolution that a qubit undergoes when decoherence is considered. The problem with this quantum channel is the fact that it is not possible to simulate it efficiently by means of classical computer as the dimension of the Hilbert state increases exponentially with the number of qubits considered [30]. This makes it impossible to construct and simulate efficient error correction codes that will be used for protecting quantum information by using conventional methods. That is why a technique named twirling is usually employed in order to obtain channels that can be managed by classical computers and that capture the essence of the GAPD channel [30, 35, 100, 101]. The significance of the twirling method comes after the fact that a correctable code for the twirled channel will also be a correctable code for the original channel [30, 102] and, thus, we can consider the simplified channels for designing codes that will eventually be successful for the actual noise. Following this logic, the most common twirling operations are the so called Pauli [30, 100, 78] and Clifford twirl approximations [30, 35, 101, 78] (PTA and CTA), where the quantum channel in consideration is averaged uniformly with the elements of the Pauli group, $\mathcal{P}$, and the elements of the Clifford group, $\mathcal{C}$, respectively. Twirling the GAPD channel with these two groups results in Pauli channels, i.e. those with Kraus operators $\{\sqrt{(1-p_{x}-p_{y}-p_{z})}\mathrm{I},\sqrt{p_{x}}\mathrm{X},\sqrt{p_{y}}\mathrm{Y},\sqrt{p_{z}}\mathrm{Z}\}$ with probabilities:
@@
The usefulness of Pauli channels resides in the fact that they can be efficiently simulated in classical computers since they fulfill the Gottesman-Knill theorem [20, 103]. Thus, we can use them in order to construct and simulate quantum error correction codes that will then be useful to protect qubits from the more general GAPD (or APD) noise [30]. Note that the CTA results in a depolarizing or symmetric Pauli channel where all the errors are equiprobable, while the PTA presents a probability bias respect to the $\mathrm{Z}$ types of errors. Therefore, the PTA is usually referred as the biased noise model, where the bias [^14] The bias is usually defined using the so called Ramsey dephasing time, $T_{2}$, that includes the dephasing induced by relaxation [35]. In this sense, the parameters relate as $1/T_{2}=1/(2T_{1})+1/T_{\phi}$ [35, 34]. is defined as $\eta=p_{z}/(p_{x}+p_{y})\approx T_{1}/T_{2}-1/2$ [104]. The bias of the channel varies significantly as a function of the technology or even as a function of the qubit of a processor being considered [30].
@@ 4.3 Noise models for multiple qubits
    where $p_{\mathrm{A}_{j}}$ is described by the PTA (biased) or CTA (depolarizing) approximations given before, and where $\mu_{T_{1}}$ and $\mu_{T_{2}}$ refer to the mean values of the relaxation and dephasing times. Taking the mean value is the usual approach.
@@
    where $p_{\mathrm{A}_{j}}(T_{1}^{j},T_{2}^{j})$ is again given by the PTA (biased) and CTA (depolarizing) approximations, but now each of the terms have particular values of relaxation and dephasing times.
@@ 4.4 SPAM and gate errors
As stated before, physical qubits experience errors from sources other than the unavoidable decoherence. Those errors refer to imperfect implementation of the operations that are done whenever the physical qubits are prepared, measured or manipulated by means of quantum gates, and are usually referred as circuit-level noise [20, 43, 107]. These errors are usually classified and modelled in the following way:
@@ 4.5 Erasure errors
In this context, an erasure channel describes the fact that a qubit at a known location has been lost with probability $p_{e}$ [110]. The fact that it is known which of the qubits is lost is important since it provides with useful information for treating those errors. The detection of leakage events in physical qubits can be done by means of the quantum jump technique or by means of ancillary qubits, for example [110, 111, 113, 114, 115]. Errors of this type with unknown locations are named deletion errors in the literature [116]. The significant difference between deletion and erasure errors lies in the fact that a deletion error leads to a decrease in the number of qubits of the system, i.e. the qubit is effectively lost, while an erasure error occurrence does not decrease the number of qubits. In this sense, the erasure channel on a qubit, $\rho$, may be described as
@@
5 Decoders for the surface code
@@
Therefore, decoding methods for error correction codes are a critical element of the code itself since their efficiency on making correct guesses of channel error will be what will determine if the method is successful or not. In this sense, the threshold of a code is a function of the decoder in question, i.e. the code can perform better or worse as a function of the method used. Making the decoder to be more accurate usually comes with the drawback of increasing its computational complexity, which ultimately makes it to be slower in making guesses. Decoders must be fast enough since the action of decoherence will not stop while estimating the error after measurement, implying that the qubit may suffer from additional errors to which the decoder will be oblivious. Thus, a slow decoder will ultimately have a bad performance (See appendix LABEL:appBack) for more details on why real time decoding is required). To sum up, the trade-off between the accuracy and complexity of the methods is vital for the field of quantum error correction [56, 57].
Surface codes can be decoded by using many methods [32, 117, 118, 119, 120, 109, 121, 46, 122, 123, 124, 125]. In this section we will explain the operation and performance, in terms of correction ability and complexity, of the main decoders for surface codes: the minimum-weight perfect matching [32, 117, 118, 119, 120], the Union-Find decoder [109], the Belief Propagation decoder [54, 46, 122] and the Tensor Network or Matrix Product State decoder [121]. In addition, we will also discuss variants of those decoding methods that have proven to be more efficient in terms of error correction ability or complexity as the Belif-Propagation Ordered Statistics Decoder (BPOSD) [46, 122], for example. Furthermore, we discuss other decoding algorithms in the literature that can be used for decoding the surface code. Many of those were proposed for decoding other topological codes such as the toric or color codes, but could, in principle, be applied for the rotated planar code. Specifically, we discuss Cellular-automaton [124], renormalization group [123], neural network or machine learning based [125] and MaxSAT [126] decoders. The end of the section includes an overview of the available software implementations available to the general public of all those decoding methods.
@@ 5.1 The Minimum Weight Perfect Matching Decoder
Before the operation of the MWPM decoder is described, some definitions of graph theory must be provided [127]. Consider a weighted graph $G$ composed by ($V$,$E$,$W$), where $V=\{v_{i}\}$ are the vertices, $E=\{e_{ij}\}$ is the set of edges which satisfy $i\neq j$ and $e_{ij}=\{v_{i},v_{j}\}$, meaning that the edge $e_{ij}$ connects the nodes $v_{i}$ and $v_{j}$, and $W=\{w_{e}\}$, $e\in E$, which is the set of weights attributed to each edge. A matching of graph $G$ is a subset of edges, denoted as $M\subseteq E$, such that for any two edges $e$ and $f$ in $M$, $e$ and $f$ do not share any common vertices. In other words, $M$ is a set of edges without common endpoints. A perfect matching is a matching that additionally satisfies the condition that every vertex in $V$ is incident to exactly one edge in $M$. A minimum weight perfect matching is the perfect matching with the smallest possible weight among all possible perfect matchings [117, 127], where the weight of a matching is defined as the sum of the weights of its edges: $\sum_{e\in M}w_{e}$. Additionally, a complete graph is a graph with the property that $\forall v_{i},v_{j}\in V,i\neq j,\exists e_{ij}\in E$.
@@
Figure 5: Graphical representation of the effect of a Pauli error in a 5x5 rotated surface code. The red lines connecting non-trivial checks represent chains.
As explained before, when the data qubits of the surface code experience a Pauli error, the checks adjacent to an odd number of errors turn into non-trivial syndrome elements. On the other hand, checks adjacent to an even number of errors will not be triggered by them, since the product of two errors of the same type will cancel. Therefore, combinations of those two types of events can be viewed as error chains over the lattice of data qubits that terminate in non-trivial checks. An example of these events is shown in FIG.5. In this sense, due to the CSS structure of the surface codes considered, two separate subgraphs where the nodes represent the non-trivial syndrome elements can be constructed: one for the $X$-checks and one for the $Z$-checks. The aforementioned nodes must be connected with all the other nodes resulting in complete subgraphs, each edge connecting two nodes must be of weight equivalent to the minimum distance between the respective checks in terms of data qubits. Once the two complete subgraphs are obtained, a perfect matching with minimum weight is seeked on those subgraphs so that the error chains corrupting the data qubits may be identified [32, 127]. One condition in order to construct a graph in which a minimum weight perfect matching can be found is that that all correctable error chains within the code have two non-trivial endpoint checks [127]. Nevertheless, the data qubits on the boundary of the code are not covered through four checks and so they can derive in chains with only one endpoint. Thus, it is necessary to consider virtual checks adjacent to the aforementioned boundary data qubits. Specifically, each non-trivial syndrome will have an associated virtual check outside the boundary of the code. Those virtual check nodes will also be connected among the other virtual checks, but the weight of those edges will be considered to be zero [127].
Figure 6: Graphical representation of a MWPM process for a specific syndrome in a 5x5 surface code.
@@
In FIG. 6, we provide an example of the operation of the MWPM decoder for a specific detected error pattern syndrome. From top to bottom and then from the left column to the right column, the first row shows the considered syndrome, where the exclamation marks correspond to the checks that have measured a non-trivial syndrome element. The second row depicts two separate graphs, one consisting of all $X$-checks (green nodes) and the other of all $Z$-checks (orange nodes). Note that in both graphs, the previously discussed virtual checks are located in the boundaries: on the left and right boundaries for the $X$-checks and on the top and bottom for the $Z$-checks. The data qubits of the code are then considered to be the edges of such graphs connecting the checks nodes with their nearest ones. Over these graphs, all the shortest paths connecting all non-trivial checks with each other and with their nearest virtual qubit are considered for the matching problem. These paths are represented with cyan, blue and violet colors for paths of weights $1$, 22 or $3$ respectively. By means of those paths, the subgraphs shown in the third row are constructed, where the weights of the edges are given by the number of data qubits that are crossed when following the path from non-trivial check to non-trivial check, i.e. each of the edges of the graphs of the second row of the figure has weight 1. Then, in the right column, the minimum weight perfect matching of those subgraphs is computed. The result of said process can be seen in the first row, notice how virtual qubits can be left unmatched. The matchings of each of the subgraphs refer to $X$ or $Z$-operators applied to the data qubits that the matching crosses. Thus, for the example in consideration, the second row of the right columns presents the operators that will be applied on the data qubits of the surface code. The $X$-checks recover $X$-errors and the $Z$-checks recover $Z$-errors. In the last row of the figure, we present the final recovery operator, where it can be seen that whenever an $X$ and $Z$-error are estimated from each of the graphs for the same data qubit, the resulting recovery operation is a $Y$-operator.
@@
The MWPM decoder estimates the Pauli error with minimum $X$ and $Z$ weight (since it considers said errors independently) that corresponds to a given syndrome. In this sense, this decoder always outputs an error whose syndrome is the same as the one measured [^16] Other decoders such as the belif-propagation decoder do not always estimate an error whose syndrome matches the true syndrome [54, 128].. In addition to always matching the syndrome, if it succeeds in matching the endpoints of an error chain to the observed syndrome it will always yield the correct outcome, even if the recovered error differs from the input one. The reason for this to happen is that the elements of the stabilizer in a 2D surface code correspond to Pauli sequences that form a closed loop in the surface code [32, 104, 129]. Thus, if the estimated error forms a closed chain with the true error occurred on the surface code, the resulting Pauli element will belong to the stabilizer set of the code and, thus, the correction will have been successful (recall eq.(6) and degeneracy of errors). In FIG. 7, we pictorically present this scenario where two error chains with same endpoints are separated by a stabilizer element. Note also that whenever an error that forms an error chain that is a loop on the data qubits of the code, all the checks will have trivial values but the codestate will not be affected by it as it will be a stabilizer element. Thus, those types of chains will be non-detectable but unharmful for the code.

### v5

@@ 4.2 Twirled quantum channels
Therefore, the introduced channel describes the evolution that a qubit undergoes when the considered decoherence processes are considered under the Markovian approximation. The problem with this quantum channel is the fact that it is not possible to simulate it efficiently by means of classical computer as the dimension of the Hilbert state increases exponentially with the number of qubits considered [27]. This makes it impossible to construct and simulate efficient error correction codes that will be used for protecting quantum information by using conventional methods. That is why a technique named twirling is usually employed in order to obtain channels that can be managed by classical computers and that capture the essence of more general channels [27, 29, 79, 80]. The significance of the twirling method comes after the fact that a correctable [^18] Correctable codes refers to the fact that the Knill-Lafalamme conditions for quantum error correction are fulfilled [59]. code for the twirled channel will also be a correctable code for the original channel [27, 81] and, thus, we can consider the simplified channels for designing codes that will eventually be successful for the actual noise. Following this logic, the most common twirling operations are the so called Pauli [27, 79, 58] and Clifford twirl approximations [27, 29, 58, 80], where the quantum channel in consideration is averaged uniformly with the elements of the Pauli group, $\mathcal{P}$, and the elements of the Clifford group, $\mathcal{C}$, respectively. Twirling the generalized amplitude and phase damping channel with these two groups results in Pauli channels, i.e. those with Kraus operators $\{\sqrt{(1-p_{x}-p_{y}-p_{z})}\mathrm{I},\sqrt{p_{x}}\mathrm{X},\sqrt{p_{y}}\mathrm{Y},\sqrt{p_{z}}\mathrm{Z}\}$ with probabilities:
@@
The usefulness of Pauli channels resides in the fact that they can be efficiently simulated in classical computers since they fulfill the Gottesman-Knill theorem [15, 82]. Thus, we can use them in order to construct and simulate quantum error correction codes that will then be useful to protect qubits from the more general noise [27]. Note that the Clifford twirl approximation results in a depolarizing or symmetric Pauli channel where all the errors are equiprobable, while the Pauli twirl approximation presents a probability bias respect to the $\mathrm{Z}$ types of errors. Therefore, the latter is usually referred as the biased noise model, where the bias [^19] The bias is usually defined using the so called Ramsey dephasing time, $T_{2}$, that includes the dephasing induced by relaxation [29]. In this sense, the parameters relate as $1/T_{2}=1/(2T_{1})+1/T_{\phi}$ [29, 28]. is defined as $\eta=p_{z}/(p_{x}+p_{y})\approx T_{1}/T_{2}-1/2$ [83]. The bias of the channel varies significantly as a function of the technology or even as a function of the qubit of a processor being considered [27].
@@ 4.3 Noise models for multiple qubits
    where $p_{\mathrm{A}_{j}}$ is described by the Pauli (biased) or Clifford (depolarizing) twirl approximations given before, and where $\mu_{T_{1}}$ and $\mu_{T_{2}}$ refer to the mean values of the relaxation and dephasing times. Taking the mean value is the usual approach.
@@
    where $p_{\mathrm{A}_{j}}(T_{1}^{j},T_{2}^{j})$ is again given by the Pauli (biased) or Clifford (depolarizing) twirl approximations, but now each of the terms have particular values of relaxation and dephasing times.
@@ 4.4 SPAM and gate errors
These refer to errors arising from the imperfect implementation of the physical operations that are done whenever the qubits are prepared, measured or manipulated by means of quantum gates and the coupling to the environment while executing them. They are usually referred to as circuit-level noise [15, 34, 87]. If only imperfect measurements are considered, the model is named phenomenological noise. These errors are usually classified and modelled in the following way:
@@ 4.5 Erasure errors
In this context, an erasure channel describes the fact that a qubit at a known location has been lost with probability $p_{e}$ [90]. The fact that it is known which of the qubits is lost is important since it provides with useful information for treating those errors. The detection of leakage events in physical qubits can be done by means of the quantum jump technique or by means of ancillary qubits, for example [90, 91, 93]. Errors of this type with unknown locations are named deletion errors in the literature [94]. The significant difference between deletion and erasure errors lies in the fact that a deletion error leads to a decrease in the number of qubits of the system, i.e. the qubit is effectively lost, while an erasure error occurrence does not decrease the number of qubits. In this sense, the erasure channel on a qubit, $\rho$, may be described as
@@
5 Decoders for surface codes
@@
Therefore, decoding methods for error correction codes are a critical element of the code itself since their efficiency on making correct guesses of channel error will be what will determine if the method is successful or not. In this sense, the threshold of a code is a function of the decoder in question, i.e. the code can perform better or worse as a function of the method used. Making the decoder to be more accurate usually comes with the drawback of increasing its computational complexity, which ultimately makes it to be slower in making guesses. Decoders must be fast enough since the action of decoherence will not stop while estimating the error after measurement, implying that the qubit may suffer from additional errors to which the decoder will be oblivious. Thus, a slow decoder will ultimately have a bad performance (See appendix A for more details on why real time decoding is required). To sum up, the trade-off between the accuracy and complexity of the methods is vital for the field of quantum error correction [43, 44].
Surface codes can be decoded by using many methods [25, 37, 60, 89, 95, 96, 97, 98, 99, 100, 101]. In this section we will explain the operation and performance, in terms of correction ability and complexity, of the main decoders for surface codes: the minimum-weight perfect matching [25, 95, 96, 97], the Union-Find decoder [89], the Belief Propagation decoder [37, 41, 98] and the Tensor Network or Matrix Product State decoder [60]. In addition, we will also discuss variants of those decoding methods that have proven to be more efficient in terms of error correction ability or complexity as the Belif-Propagation Ordered Statistics Decoder (BPOSD) [37, 98], for example. Furthermore, we discuss other decoding algorithms in the literature that can be used for decoding different types of surface codes. Many of those were proposed for decoding topological codes such as the toric or color codes, but could, in principle, be applied for the rotated planar code. Specifically, we discuss cellular-automaton [100], renormalization group [99], neural network or machine learning based [101] and MaxSAT [102] decoders. The end of the section includes an overview of the available software implementations available to the general public of all those decoding methods.
@@ 5.1 The Minimum Weight Perfect Matching Decoder
Before the operation of the Minimum Weight Perfect Matching decoder is described, some definitions of graph theory must be provided [103]. Consider a weighted graph $G$ composed by ($V_{G}$$E_{G}$$W_{G}$), where $V_{G}=\{v_{i}\}$ are the vertices, $E_{G}=\{e_{ij}\}$ is the set of edges which satisfy $i\neq j$ and $e_{ij}=\{v_{i},v_{j}\}$, meaning that the edge $e_{ij}$ connects the nodes $v_{i}$ and $v_{j}$, and $W_{G}=\{w_{e}\}$, $e\in E$, which is the set of weights attributed to each edge. A matching of graph $G$ is a subset of edges, denoted as $M\subseteq E_{G}$, such that for any two edges $e$ and $f$ in $M$, $e$ and $f$ do not share any common vertices. In other words, $M$ is a set of edges without common endpoints. A perfect matching is a matching that additionally satisfies the condition that every vertex in $V$ is incident to exactly one edge in $M$. A minimum weight perfect matching is the perfect matching with the smallest possible weight among all possible perfect matchings [95, 103], where the weight of a matching is defined as the sum of the weights of its edges: $\sum_{e\in M}w_{e}$. Additionally, a complete graph is a graph with the property that $\forall v_{i},v_{j}\in V,i\neq j,\exists e_{ij}\in E_{G}$. The Minimum Weight Perfect Matching (MWPM) decoder makes use of these concepts from graph theory in order to develop an algorithm for decoding.
@@
Figure 5: Graphical representation of the effect of a Pauli error in a 5x5 rotated surface code. $Z$-errors trigger orange $Z$-checks and $X$-errors trigger green $X$-checks. The red lines connecting non-trivial checks represent chains (when two data qubits trigger the same check it becomes trivial).
As explained before, when the data qubits of the surface code experience a non-trivial Pauli error, the checks adjacent to an odd number of errors turn into non-trivial syndrome elements while checks adjacent to an even number of errors are not triggered by them, since the product of two errors of the same type will cancel. If we are to consider $X$ and $Z$-errors separately [^21] Recall that $Y$-errors can be regarded as combinations of both $X$ and $Z$-errors on the same qubit., we can regard the error which affected the code as a set of chains of data qubits which have non-trivial checks on their endpoints. An example of this consideration is shown in FIG. 5, notice how, adjacent $X$-errors on the bottom-left and bottom-right side of the figure form a chain denoted by a red line with non-trivial checks at its endpoints. There can also be the chance of errors ending in a boundary where there are no checks susceptible to them, this is the example of the top $Z$-error in FIG. 5. For the sake of the MWPM decoding method, we are interested in chains having non-trivial checks on their endpoints, therefore, we consider the existence of virtual checks, which are hypothetical XX($Z$) check qubits on the left and right (top and bottom) boundaries that can be considered to be either trivial or non-trivial. Moving back to the top $Z$-error from FIG 5, the top right non-trivial lighter color $Z$-check is a virtual check, since there is no physical check in that spatial position, we consider it so as to be able to have two non-trivial checks at the endpoints on the $Z$-chain.
In this sense, due to the CSS structure of the rotated planar code, two separate graphs can be constructed: one for the $X$-checks and the other for the $Z$-checks, both combined are known as the syndrome graph [104]. Within the syndome graph, nodes represent the checks and virtual checks, either $X$-checks in the $X$-subgraph or $Z$-checks in the $Z$-subgraph. Also, edges represent data qubits, that is, if two checks are susceptible to a change in the parity of a specific data qubit ($X$-parity for the $X$-checks and $Z$-parity for the $Z$-checks), the two nodes which represent said checks will be connected by an edge representing that specific data qubit. This could be the case of the two non-trivial $X$-checks on the bottom-left side of FIG. 5, which are both susceptible to the change in $X$-parity of the data qubit in between them, thus, within the constructed $X$-subgraph, there is an edge in between them. All edges in the syndrome graph have the same weight (commonly 1), since we consider all data qubits to undergo non-trivial Pauli operators with the same probability.
Given a specific syndrome, we want to consider the error chains which may have produced it. As discussed earlier, chains have non-trivial checks on their endpoints, therefore, we can study this problem by looking for all possible same type non-trivial check pairs. Afterwards, we can create the decoding graph [104]. In the decoding graph, nodes correspond to non-trivial syndrome checks and some virtual checks. Moreover, it can also be separated into two independent subgraphs, the $X$-decoding subgraph and the $Z$-one. Nodes correspondent to non-trivial checks within each of the subgraphs will be connected with edges to all the other non-trivial checks of the same type, altogether with their nearest virtual-check. For example, in FIG. 5, the nearest virtual check of the non-trivial $Z$-check in the bulk of the code will be the aforementioned top-right virtual check. Additionally, the weight of each edge in the decoding graph will be given by the weight of the minimum distance path in between both checks within the syndrome graph. For instance, going back to FIG. 5, if we were to construct the $Z$-decoding subgraph, we would observe that is only composed by a node representing the non-trivial $Z$-check close to the center of the code and the virtual check from the top-right, connected via an edge of weight 2.
Once the two decoding subgraphs are obtained, one can compute their minimum weight perfect matching [25, 103]. The minimum weight perfect matching will be the chosen set of chains which is considered to have produced the syndrome. Afterwards, the edges belonging to said set of chains are mapped back to the syndrome graph, and the data qubits whose edges belong to the minimum weight perfect matching are considered to have undergone the non-trivial Pauli operators from the error. The decoding $X$-subgraph will return $X$-errors and the decoding $Z$-subgraph will return $Z$-errors.
Figure 6: Graphical representation of a MWPM decoding process for a specific syndrome in a 5x5 surface code. In a) there is a representation of a syndrome of a distance-5 rotated code. Images from b) represent the syndrome graphs of the distance-5 rotated code splitted in the $X$ and $Z$-check graphs and the illustrations from c) represent the decoding graphs for such syndrome for each check type. d) represents the obtained matchings from the decoding graphs and e) the recovered error for every subgraph within the syndrome graph. Finally f) shows the recovered error.
@@
In FIG. 6, we provide an example of the operation of the MWPM decoder for a specific detected error pattern syndrome. FIG. 6 a) shows the considered syndrome, where the exclamation marks correspond to the checks that have measured a non-trivial syndrome element. FIG. 6 b) depicts the two syndrome subgraphs, the $X$-subgraph (green nodes) and the $Z$-subgraph (orange nodes), edges in the $X$ and $Z$-subgraphs are denoted with green and orange lines respectively. Note that in both graphs, the previously discussed virtual checks are located in the boundaries: on the left and right boundaries for the $X$-checks and on the top and bottom for the $Z$-checks. The edges connecting nodes in the syndrome graph represent the data qubits which, when changing parity, trigger a change in the parity of the check nodes they are. Over these graphs, all the shortest paths connecting all non-trivial checks with each other and with their nearest virtual qubit are considered for the matching problem. These paths are represented with cyan, blue and violet colors for paths of weights $1$, 22 or $3$ respectively. By means of those paths, the decoding subgraphs shown in FIG. 6 c) are constructed, where the weights of the edges are given by the number of data qubits that are crossed when following the path from non-trivial check to non-trivial check, i.e. each of the edges of the graphs of the second row of the figure has weight 1. Then, in FIG. 6 d), the minimum weight perfect matching of those subgraphs is computed. The result of said process can be seen in FIG. 6 e), notice how virtual qubits can be left unmatched. The matchings of each of the subgraphs refer to $X$ or $Z$ operators applied to the data qubits that the matching crosses. Thus, for the example in consideration, the second row of the right columns presents the operators that will be applied on the data qubits of the rotated code. The $X$-checks recover $X$-errors and the $Z$-checks recover $Z$-errors. In FIG. 6 f), we present the final recovery operator, where it can be seen that whenever an $X$ and $Z$-error are estimated from each of the graphs for the same data qubit, the resulting recovery operation is a $Y$ operator.
@@
The MWPM decoder estimates the Pauli error with minimum $X$ and $Z$ weight (since it considers said errors independently) that corresponds to a given syndrome. In this sense, this decoder always outputs an error whose syndrome is the same as the one measured [^22] Other decoders such as the belief propagation decoder do not always estimate an error whose syndrome matches the true syndrome [41, 105].. In addition to always matching the syndrome, if the recovered minimum weight perfect matching from the decoding graph succeeds, the recovered error will be correct independently of the path chosen within the syndrome graph. In other words, if the error can be represented as a set of chains which relates the same non-trivial checks as the minimum weight perfect matching, the recovered error will correct it even if they are not the same Pauli operator. The reason for this to happen is that the elements of the stabilizer in a 2D surface code correspond to Pauli sequences that form a closed loop in the surface code [25, 83]. Thus, if the estimated error forms a closed chain with the true error occurred on the surface code, the resulting Pauli element will belong to the stabilizer set of the code and, thus, the correction will have been successful (recall eq.(6) and degeneracy of errors). In FIG. 7, we pictorically present this scenario where two error chains with same endpoints are separated by a stabilizer element. Note also that whenever an error that forms an error chain that is a loop on the data qubits of the code, all the checks will have trivial values but the codestate will not be affected by it as it will be a stabilizer element. Thus, those types of chains will be non-detectable but unharmful for the code.